Can I get it solved step by step please

Consider a 30-year U.S. municipal bond with 25 years left to maturity. The bond carries an 6.5 percent coupon and is priced at $985? What is the after-tax yield to maturity for an investor with a 30 percent marginal tax rate?
Answer #1

Assuming face value to be $1,000

Coupon = 0.065 * 1000 = 65

Yield to maturity = 6.6244%

Keys to use in a financial calculator: FV 1000, PV -985, N 25, PMT 65, CPT I/Y

after tax yield to maturity = 0.066244 (1 - 0.3)

after tax yield to maturity = 0.066244 * 0.7

after tax yield to maturity = 4.64%